Ezekiel333:

After a devastating winter, when thousands of fish died, an environmental scientist has replenished the trout stock in a fishing pond. He started with 5,000 baby trout and has finished a count to find that, in 8 years, the population is estimated to be 33,000. Assuming an exponential growth pattern, what is the annual growth rate (rounded to the nearest tenth of a percent) of the new trout population? Hint: A(t) = A0(1 + r)t, where A(t) is the final amount, A0 is the initial amount, r is the growth rate expressed as a decimal, and t is time. (4 points) 26.6% 2.7% 79.0%

Vocaloid:

The problem states that the growth can be modeled with the equation A(t) = A0(1 + r)^t Where A0 is the initial population (5,000), A(t) is the final population (33,000), t is time (8 yrs) and r is the growth rate. Plug in the appropriate quantities and solve for r.

Ezekiel333:

33000 = 5000 ( 1 + r ) ^8

Vocaloid:

Good keep going

Vocaloid:

I usually just use a calculator like wolfram alpha but if you want to do this by hand, divide both sides by 5000, take the log8 of both sides to cancel out that 8 exponent, then subtract 1 from both sides

Ezekiel333:

What would the set up be

Vocaloid:

You could just type the equation into a calculator and have it solve for r

Ezekiel333:

wolfram is being slow

Vocaloid:

Mathway maybe?

Ezekiel333:

1 + 8 r + 28 r 2 + 56 r 3 + 70 r 4 + 56 r 5 + 28 r 6 + 8 r 7 + r 8

Ezekiel333:

thats what i got from mathway it is showing a rate that is linear

Ezekiel333:

the numbers after the "r" are the exponents

Vocaloid:

If you type 33000 = 5000(1+x)^8 it’ll solve for x to get x = 0.266 or 26.6% growth

Ezekiel333:

hm smart

Ezekiel333:

still the same thing

Ezekiel333:

i tried every solution

Ezekiel333:

26.6% 2.7% 79.0% has to be one of these

Vocaloid:

The solution to 33000 = 5000(1+x)^8 is x = 0.266 or 26.6% growth

Ezekiel333:

ohhh